OpenStudy (anonymous):

solve: log base 5 (x-9)=.833

OpenStudy (slaaibak):

\[\log_5 (x-9) = {5 \over 6}\] so \[5^{5 \over 6} = x - 9\] \[x = 5^{5 \over 6} + 9 = 12.82362246\]

OpenStudy (anonymous):

thanks!

OpenStudy (slaaibak):

cool